OpenStudy (anonymous):

8 1/2 divided by 2 1/4

OpenStudy (anonymous):

4 1/8

OpenStudy (anonymous):

or 33/8

OpenStudy (anonymous):

@slade @.Gjallarhorn. both of those are not options and this time i did type it right

OpenStudy (anonymous):

am i right @audieeeuniverse

OpenStudy (kash_thesmartguy):

Give us the options!

OpenStudy (anonymous):

OK HERE = 17/2 ÷ 9/4 = (17 × 4) / (9 × 2) = 68/18 = 34/9 = 3 7/9 < answer????

OpenStudy (anonymous):

a) 4 1/2 b) 3 1/2 c) 3 7/9

OpenStudy (anonymous):

A

OpenStudy (anonymous):

so C according to my answer

OpenStudy (anonymous):

no its not @Kash_TheSmartGuy tell them it A

OpenStudy (study_together):

\[\frac{ 17 }{ 2 }\div \frac{ 9 }{ 4 }\]\[\frac{ 17 }{ 2 }\times \frac{ 4 }{ 9 }\]\[\frac{ 34 }{ 9 }\]\[3\frac{ 7 }{ 9 }\]\[3\frac{ 7 }{ 9 }= Answer\]

OpenStudy (anonymous):

http://www.calculatorsoup.com/

OpenStudy (anonymous):

ill proove it

OpenStudy (anonymous):

THAT MEDAL MADE MY SMART SCORE 50 LES GOO

OpenStudy (anonymous):

click the link its proof